# Neural Networks and (Virtual) Extended Formulations

### v2

@@ 1.2. Further Related Work
In terms of exact representation, it is known that a function can be represented if and only if it is CPWL (Arora et al., 2018), and it is still an open question whether constant depth is sufficient to do so (Hertrich et al., 2023; Haase et al., 2023). Interestingly, also for this question, monotone networks seem to be more amenable for proving lower bounds than their non-monotone counterparts (Valerdi, 2024). Furthermore, the related question of how to efficiently write a non-convex CPWL function as a difference of two convex ones received quite some attention recently (Brandenburg et al., 2024a; Tran and Wang, 2024).
@@ 2. Preliminaries
For a polytope $P\subseteq\mathbb{R}^{d}$, its support function $f_{P}\colon\mathbb{R}^{d}\to\mathbb{R}$ is defined as $f_{P}(c)=\max_{x\in P}c^{\top}x$. In other words, it maps a linear objective direction to the optimal value obtained by optimizing over the polytope. Support functions are convex, continuous, piecewise linear, and positively homogeneous functions; here, a function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is positively homogeneous if $f(\lambda x)=\lambda f(x)$ for all scalars $\lambda\geq 0$. Let $\mathcal{F}^{d}$ be the set of all such functions from $\mathbb{R}^{d}\to\mathbb{R}$ and let $\mathcal{P}^{d}$ be the set of all polytopes embedded in $\mathbb{R}^{d}$. Then the map $\mathcal{P}^{d}\to\mathcal{F}^{d}$, $P\mapsto f_{P}$, is a bijection that is compatible with a certain set of operations:
@@
The inverse map that maps a support function $f_{P}$ to its unique associated polytope $P$ is well-studied in tropical geometry. In fact, borrowing the name from (tropical) polynomials, we call $P$ the Newton polytope of $f_{P}$. The connection between CPWL functions and their Newton polytopes has previously been used in the study of neural networks and their expressivity (Zhang et al., 2018; Montúfar et al., 2022; Haase et al., 2023; Hertrich et al., 2023; Brandenburg et al., 2024b; Misiakos et al., 2022).
@@
We sometimes focus on rank-2 maxout networks. To justify this, note that a rank-$k$ maxout unit can be simulated with $k-1$ rank-2 maxout units, which allows to transfer size bounds to other ranks if $k$ is viewed as a fixed constant. Moreover, observe that a rank-2 maxout network can be simulated by three parallel ReLU units because $\max\{x,y\}=\max\{0,x-y\}+\max\{0,y\}-\max\{0,-y\}$. Note, however, that this introduces negative weights to the ReLU network even if the maxout network was monotone. Therefore, unlike in the non-monotone case, monotone ReLU networks are indeed weaker than monotone maxout networks.
@@ 3. Lower Bounds for Monotone and Input-Convex Neural Networks
The goal of this section is to prove the following theorem and use it to derive exponential lower bounds on monotone and input-convex neural networks. Here and in the following, we defer some of the proofs to Appendix A.
@@
$y_{v}\geq\sum_{u\in\delta^{\mathrm{in}}_{v}}w_{uv}^{i}y_{u}\quad\forall\ i=1,\dots,k\quad\forall\text{ maxout units }v\enspace.$
@@
$y_{v}\geq\sum_{u\in\delta^{\mathrm{in}}_{v}}w_{uv}^{i}y_{u}\quad\forall\ i=1,\dots,k\quad\Leftrightarrow\quad y_{v}\geq\max_{i=1,\dots,k}\sum_{u\in\delta^{\mathrm{in}}_{v}}w_{uv}^{i}y_{u}\enspace.$
@@ 4.1. Neural Networks are Virtual Extended Formulations
The proof of this theorem requires three steps, two of which are already available through Propositions 3.3 and 3.5. In addition, we need another proposition showing that every maxout network can be written as the difference of two monotone maxout networks each of which has the same size as the original network. This is related to writing functions represented by neural networks as tropical rational functions (Zhang et al., 2018; Brandenburg et al., 2024b).
@@
If $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is represented by a rank-$k$ maxout network of size $s$, then it can be written as a difference $f=g-h$ of two functions $g$ and $h$ that are both representable with a monotone rank-$k$ maxout network of size $s$.
@@ 4.2. Optimizing over Virtual Extended Formulations
The key idea is that we need to make sure that ties between equally good solutions are broken consistently across $Q$ and $R$. Therefore, we suggest a deterministic procedure based on a lexicographic objective function. By this, we mean a sequence of objective functions where later entries are only considered to break ties.
@@ Acknowledgments
Appendix A Omitted Proofs
Proof of Proposition 3.2.
To go from $\mathcal{N}_{p-1}$ to $\mathcal{N}_{p}$ for some $p\geq 1$, we let the first $p-1$ maxout neurons unchanged. Let $I\subseteq V$ be the set of input neurons. Consider the weights $w_{uv}^{i}$ on arcs entering $v\coloneqq v_{p}$ from an input neuron $u\in I$. Let $\gamma_{u}\coloneqq\min_{i=1,\dots,k}w_{uv}^{i}$ be the smallest such weight for each $u\in I$, which might be negative or positive. Observe that
(1) $z_{v}$ $=\max\Big{\{}\sum_{u\in I}(w_{uv}^{i}-\gamma_{u})z_{u}+\sum_{u\in\delta^{\mathrm{in}}_{v}\setminus I}w_{uv}^{i}z_{u}\Bigm{|}i=1,\dots,k\Big{\}}+\sum_{u\in I}\gamma_{u}z_{u}\enspace.$
To this end, we first show by induction on $p=1,\dots,s-1$ that for a negative unit vector $x=-e_{j}$ as input, each neuron $v=v_{p}$ outputs $z_{v}(-e_{j})=0$. If $p=1$, then the only incoming edges of $v=v_{1}$ are coming from input neurons. Moreover, since we are inputting $-e_{j}$, only the $j$-th input neuron, call it $u$, propagates a nonzero value. By the choice of $\gamma_{u}$, there must be an index $i$ such that our updated weight $w_{uv}^{i}-\gamma_{u}$ is exactly zero, while it is nonnegative for all other indices. Therefore, upon input $-e_{j}$, the maximum expression evaluates to 00, settling the induction start. For the induction step, observe that the very same argument applies, since by induction all previous neurons output zero upon input $-e_{j}$.
As a consequence of the latter claim, we obtain that the output $z_{v_{s}}(-e_{j})$ of $\mathcal{N}_{s-1}$ is precisely the negated weight from $u$ to $v_{s}$. Since $z_{v_{s}}(0)=0$, monotonicity of the final function implies that the output on $-e_{j}$ must be nonpositive, implying that the weight must be nonnegative, as claimed. Thus, we just showed that also the incoming weights of $v_{s}$ must be nonnegative in $\mathcal{N}_{s-1}$, implying that $\mathcal{N}_{s-1}$ is monotone. ∎
Proof of Proposition 3.5..
The key idea behind this statement is the well-known fact that the dual of the epigraph of $f_{P}$ as a cone in $\mathbb{R}^{d+1}$ equals the homogenization of $P$, see, e.g., Bertsekas [2009], Rockafellar [1970]. We give a short geometric proof of this fact to show the ingredients of the construction. A point $(c,t)$ in the epigraph $\operatorname{epi}(f_{P})$ fulfills the equivalent conditions
This yields the equivalent representation
$\operatorname{epi}(f_{P})=\Big{\{}(c,t)\in\mathbb{R}^{d}\times\mathbb{R}\Bigm{|}(c^{\top},t)\begin{pmatrix}x\\ -1\end{pmatrix}\leq 0\;\forall x\in P\Big{\}}\enspace.$
This is the polar cone of the cone over $P\times\{-1\}$ that is
$\operatorname{epi}(f_{P})=\text{cone}\left(\left\{(x,-1)\mid x\in P\right\}\right)^{\star}\enspace.$
To finish the proof, we argue that (de-)homogenization and cone polarity preserve extension complexity. The claim about (de-)homogenization follows from the discussion before Gouveia et al. [2013, Theorem 6]. The statement about cone polarity is implied by the discussion before Gouveia et al. [2013, Proposition 2]. ∎
Proof of Theorem 3.9..
With that modification, $Q$ is a “monotone polytope” [^4] Note that this notion of a monotone polytope does not correspond to the notion of monotone neural networks. as defined in Rothvoß [2017].
In conclusion, also for $Q$ we have the inequality $f_{\mathrm{MAT}}(c)\leq f_{Q}(c)\leq(1+\epsilon)f_{\mathrm{MAT}}(c)$ for all $c\in\mathbb{R}_{\geq 0}^{E}$. As argued around Rothvoß [2017, Equation 7], this is equivalent to $P_{\mathrm{MAT}}\subseteq Q\subseteq(1+\epsilon)P_{\mathrm{MAT}}$. By Rothvoß [2017, Corollary 4.2], this implies that $\operatorname{xc}(Q)\geq 2^{\Omega(\min\{n,1/\epsilon\})}$; see also Sinha [2018], Braun and Pokutta [2014]. The definition (2) of $Q$ implies $\operatorname{xc}(Q)\leq\operatorname{xc}(P_{f})+2\cdot\lvert E\rvert\leq\operatorname{xc}(P_{f})+\mathcal{O}(n^{2})$. Thus, we obtain the same asymptotic lower bound on $\operatorname{xc}(P_{f})$. The statement then follows by Theorem 3.1. ∎
Proof of Proposition 4.2..
To go from $\mathcal{N}_{p-1}$ to $\mathcal{N}_{p}$ for some $p\geq 1$, we let the first $p-1$ neurons unchanged. Now consider the weights $w_{uv}^{i}$ of $v\coloneqq v_{p}$. We split them into a positive and negative part, that is, $w_{uv}^{i}=a_{uv}^{i}-b_{uv}^{i}$ with $a_{uv}^{i}=\max\{w_{uv}^{i},0\}$, $b_{uv}^{i}=\max\{-w_{uv}^{i},0\}$. Observe that
(3) $z_{v}$ $=\max\Big{\{}\sum_{u\in\delta^{\mathrm{in}}_{v}}\big{(}a_{uv}^{i}+\sum_{j\neq i}b_{uv}^{j}\big{)}z_{u}\Bigm{|}i=1,\dots,k\Big{\}}-\sum_{u\in\delta^{\mathrm{in}}_{v}}\sum_{j=1}^{k}b_{uv}^{j}z_{u}\enspace.$
Proof of Proposition 4.5..
Proof of Theorem 4.6..

### v3

@@ 1.2. Further Related Work
In terms of exact representation, it is known that a function can be represented if and only if it is CPWL (Arora et al., 2018), and it is still an open question whether constant depth is sufficient to do so (Hertrich et al., 2023; Haase et al., 2023; Grillo et al., 2025; Bakaev et al., 2025b). Interestingly, also for this question, monotone networks seem to be more amenable for proving lower bounds than their non-monotone counterparts (Valerdi, 2024; Bakaev et al., 2025a). Furthermore, the related question of how to efficiently write a non-convex CPWL function as a difference of two convex ones received quite some attention recently (Brandenburg et al., 2025; Tran and Wang, 2024).
@@ 2. Preliminaries
For a polytope $P\subseteq\mathbb{R}^{d}$, its support function $f_{P}\colon\mathbb{R}^{d}\to\mathbb{R}$ is defined as $f_{P}(c)=\max_{x\in P}c^{\top}x$. In other words, it maps a linear objective direction to the optimal value obtained by optimizing over the polytope. Support functions are convex, continuous, piecewise linear, and positively homogeneous functions; and every function with these properties is a support function of a polytope. Here, a function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is positively homogeneous if $f(\lambda x)=\lambda f(x)$ for all scalars $\lambda\geq 0$. Let $\mathcal{F}^{d}$ be the set of all such functions from $\mathbb{R}^{d}\to\mathbb{R}$ and let $\mathcal{P}^{d}$ be the set of all polytopes embedded in $\mathbb{R}^{d}$. Then the map $\mathcal{P}^{d}\to\mathcal{F}^{d}$, $P\mapsto f_{P}$, is a bijection that is compatible with a certain set of operations:
@@
The inverse map that maps a support function $f_{P}$ to its unique associated polytope $P$ is well-studied in tropical geometry. In fact, borrowing the name from (tropical) polynomials, we call $P$ the Newton polytope of $f_{P}$. The connection between CPWL functions and their Newton polytopes has previously been used in the study of neural networks and their expressivity (Zhang et al., 2018; Montúfar et al., 2022; Haase et al., 2023; Hertrich et al., 2023; Brandenburg et al., 2024; Misiakos et al., 2022).
@@
We sometimes focus on rank-2 maxout networks. To justify this, note that a rank-$k$ maxout unit can be simulated with k1k-1 rank-2 maxout units, which allows to transfer size bounds to other ranks if $k$ is viewed as a fixed constant. Moreover, observe that a rank-2 maxout network can be simulated by three parallel ReLU units because $\max\{x,y\}=\max\{0,x-y\}+\max\{0,y\}-\max\{0,-y\}$. Note, however, that this introduces negative weights to the ReLU network even if the maxout network was monotone. Therefore, unlike in the non-monotone case, monotone ReLU networks are indeed weaker than monotone maxout networks.
@@ 3. Lower Bounds for Monotone and Input-Convex Neural Networks
The goal of this section is to prove the following theorem and use it to derive exponential lower bounds on monotone and input-convex neural networks.
@@
Proof.
To go from $\mathcal{N}_{p-1}$ to $\mathcal{N}_{p}$ for some $p\geq 1$, we let the first p1p-1 maxout neurons unchanged. Let $I\subseteq V$ be the set of input neurons. Consider the weights $w_{uv}^{i}$ on arcs entering $v\coloneqq v_{p}$ from an input neuron $u\in I$. Let $\gamma_{u}\coloneqq\min_{i=1,\dots,k}w_{uv}^{i}$ be the smallest such weight for each $u\in I$, which might be negative or positive. Observe that
(1) $z_{v}$ $=\max\Big\{\sum_{u\in I}(w_{uv}^{i}-\gamma_{u})z_{u}+\sum_{u\in\delta^{\mathrm{in}}_{v}\setminus I}w_{uv}^{i}z_{u}\Bigm|i=1,\dots,k\Big\}+\sum_{u\in I}\gamma_{u}z_{u}\kern 4.62497pt.$
To this end, we first show by induction on $p=1,\dots,s-1$ that for a negative unit vector $x=-e_{j}$ as input, each neuron $v=v_{p}$ outputs $z_{v}(-e_{j})=0$. If $p=1$, then the only incoming edges of $v=v_{1}$ are coming from input neurons. Moreover, since we are inputting $e_{j}$, only the $j$-th input neuron, call it $u$, propagates a nonzero value. By the choice of $\gamma_{u}$, there must be an index $i$ such that our updated weight $w_{uv}^{i}-\gamma_{u}$ is exactly zero, while it is nonnegative for all other indices. Therefore, upon input $e_{j}$, the maximum expression evaluates to 0, settling the induction start. For the induction step, observe that the very same argument applies, since by induction all previous neurons output zero upon input $e_{j}$.
As a consequence of the latter claim, we obtain that the output $z_{v_{s}}(-e_{j})$ of $\mathcal{N}_{s-1}$ is precisely the negated weight from $u$ to $v_{s}$. Since $z_{v_{s}}(0)=0$, monotonicity of the final function implies that the output on $e_{j}$ must be nonpositive, implying that the weight must be nonnegative, as claimed. Thus, we just showed that also the incoming weights of $v_{s}$ must be nonnegative in $\mathcal{N}_{s-1}$, implying that $\mathcal{N}_{s-1}$ is monotone. ∎
@@
$y_{v}\geq\sum_{u\in\delta^{\mathrm{in}}_{v}}w_{uv}^{i}y_{u}\quad\forall\ i=1,\dots,k\quad\forall\text{ maxout units }v\kern 5.0pt.$
@@
$y_{v}\geq\sum_{u\in\delta^{\mathrm{in}}_{v}}w_{uv}^{i}y_{u}\quad\forall\ i=1,\dots,k\quad\Leftrightarrow\quad y_{v}\geq\max_{i=1,\dots,k}\sum_{u\in\delta^{\mathrm{in}}_{v}}w_{uv}^{i}y_{u}\kern 5.0pt.$
@@
Proof.
The key idea behind this statement is the well-known fact that the dual of the epigraph of $f_{P}$ as a cone in $\mathbb{R}^{d+1}$ equals the homogenization of $P$, see, e.g., Bertsekas (2009); Rockafellar (1970). We give a short geometric proof of this fact to show the ingredients of the construction. A point $(c,t)$ in the epigraph $\operatorname{epi}(f_{P})$ fulfills the equivalent conditions
This yields the equivalent representation
$\operatorname{epi}(f_{P})=\Big\{(c,t)\in\mathbb{R}^{d}\times\mathbb{R}\Bigm|(c^{\top},t)\begin{pmatrix}x\\ -1\end{pmatrix}\leq 0\;\forall x\in P\Big\}\kern 5.0pt.$
This is the polar cone of the cone over $P\times\{-1\}$ that is
$\operatorname{epi}(f_{P})=\text{cone}\left(\left\{(x,-1)\mid x\in P\right\}\right)^{\star}\kern 5.0pt.$
To finish the proof, we argue that (de-)homogenization and cone polarity preserve extension complexity. The claim about (de-)homogenization follows from the discussion before Gouveia et al. (2013, Theorem 6). The statement about cone polarity is implied by the discussion before Gouveia et al. (2013, Proposition 2). ∎
@@
Proof.
With that modification, $Q$ is a “monotone polytope” [^4] Note that this notion of a monotone polytope does not correspond to the notion of monotone neural networks. as defined in Rothvoß (2017).
In conclusion, also for $Q$ we have the inequality $f_{\mathrm{MAT}}(c)\leq f_{Q}(c)\leq(1+\epsilon)f_{\mathrm{MAT}}(c)$ for all $c\in\mathbb{R}_{\geq 0}^{E}$. As argued around Rothvoß (2017, Equation 7), this is equivalent to $P_{\mathrm{MAT}}\subseteq Q\subseteq(1+\epsilon)P_{\mathrm{MAT}}$. By Rothvoß (2017, Corollary 4.2), this implies that $\operatorname{xc}(Q)\geq 2^{\Omega(\min\{n,1/\epsilon\})}$; see also Sinha (2018); Braun and Pokutta (2014). The definition (2) of $Q$ implies $\operatorname{xc}(Q)\leq\operatorname{xc}(P_{f})+2\cdot\lvert E\rvert\leq\operatorname{xc}(P_{f})+\mathcal{O}(n^{2})$. Thus, we obtain the same asymptotic lower bound on $\operatorname{xc}(P_{f})$. The statement then follows by Theorem 3.1. ∎
@@ 4.1. Neural Networks are Virtual Extended Formulations
The proof of this theorem requires three steps, two of which are already available through Propositions 3.3 and 3.5. In addition, we need another proposition showing that every maxout network can be written as the difference of two monotone maxout networks each of which has the same size as the original network. This is related to writing functions represented by neural networks as tropical rational functions (Zhang et al., 2018; Brandenburg et al., 2024).
@@
If $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is represented by a rank-$k$ maxout network of size $s$, then it can be written as a difference f=ghf=g-h of two functions $g$ and $h$ that are both representable with a monotone rank-$k$ maxout network of size $s$.
Proof.
To go from $\mathcal{N}_{p-1}$ to $\mathcal{N}_{p}$ for some $p\geq 1$, we let the first p1p-1 neurons unchanged. Now consider the weights $w_{uv}^{i}$ of $v\coloneqq v_{p}$. We split them into a positive and negative part, that is, $w_{uv}^{i}=a_{uv}^{i}-b_{uv}^{i}$ with $a_{uv}^{i}=\max\{w_{uv}^{i},0\}$, $b_{uv}^{i}=\max\{-w_{uv}^{i},0\}$. Observe that
(3) $z_{v}$ $=\max\Big\{\sum_{u\in\delta^{\mathrm{in}}_{v}}\big(a_{uv}^{i}+\sum_{j\neq i}b_{uv}^{j}\big)z_{u}\Bigm|i=1,\dots,k\Big\}-\sum_{u\in\delta^{\mathrm{in}}_{v}}\sum_{j=1}^{k}b_{uv}^{j}z_{u}\kern 4.62497pt.$
@@ 4.2. Optimizing over Virtual Extended Formulations
The key idea is that we need to make sure that ties between equally good solutions are broken consistently across $Q$ and $R$. Therefore, we suggest a deterministic procedure based on a lexicographic objective function, which is a standard notion in multi-criteria optimization. By this, we mean a sequence of objective functions where later entries are only considered to break ties.